OpenStudy (anonymous):

Simplify the expression: cot x sin x - sin((pi/2)-x) + cos x.

OpenStudy (anonymous):

\[\cot x \sin x - \sin \left( \frac{ \pi }{ 2 }-x \right) + \cos x\]

OpenStudy (anonymous):

I will draw what I did hang on.

OpenStudy (ali2x2):

cos x - cos x + cos x= cos x

OpenStudy (anonymous):

\[\frac{ \cos x }{ \sin x }*\sin x - \cos x +\cos x\]

OpenStudy (ali2x2):

:P

OpenStudy (anonymous):

Then I was left with cos x.

OpenStudy (anonymous):